Tohban Report 2016-05-25
Tohban Reports | |
---|---|
Start Date: | 18 May 2016 |
End Date: | 25 May 2016 |
Tohban: | Nicole Duncan |
Tohban email: | nicoleduncan@berkeley.edu |
Next Tohban: | Jim |
List all reports |
Next Meeting
For SPD next week we are changing the meeting time to 2pm PDT, 3pm Mtn (SPD) and 5pm EST.
Solar Activity
The activity has been very low. There are currently 2 ARs on disk with very little complexity. Expecting very quiet Sun in the next week.
How many GOES flares occurred?
Flares above B, C, M, X class were 21 4 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 4 2 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 3 2 0 0
There were RHESSI flares/GOES flares 87 / 25 over the time range 16-May-16 23-May-16
Memory Management
Decimation was in normal/vigorous for the full time period. The SSR is empty after each day's passes, with a maximum of 12 %.
Spacecraft Status
The cold plates 1 and 2 are at 136.0 and 134.5 K, down a few tenths from last week.
We are 2 W over the linear interpolation of the cryocooler power curve.
D1 and 9 were turned off on May 18. D3 and D8 are on and continue to record events.
Data Gaps
DATA GAPS FOR 2016/05/20 TIME RANGE: 2016-05-20T00:00:00.000 -- 2016-05-21T00:00:00.000 GAPS IN APP_ID = 1 (VC1-SOH) WITH PACKET RATE LT 30 N_GAPS 2 GAP START TIME GAP END TIME GAP (SEC) 2016-05-20T18:25:00.000 -- 2016-05-20T18:35:00.000 600.00000 2016-05-20T19:40:00.000 -- 2016-05-20T19:45:00.000 300.00000 NO GAPS IN APP_ID = 154 (VC1-PMTRAS) GAPS IN APP_ID = 102 (VC3-MONITOR RATES) WITH PACKET RATE LT 30 N_GAPS 2 GAP START TIME GAP END TIME GAP (SEC) 2016-05-20T19:30:00.000 -- 2016-05-20T19:35:00.000 300.00000 2016-05-20T21:20:00.000 -- 2016-05-20T22:45:00.000 5100.0000
Detector issues
D8 has shown evidence of the light leak issue, whereas it previously wasn't present in this detector before this anneal. There was discussion as to if this correlation between D8 & D3 was seen prior to the 2014 anneal as well. Old Tohban reports should be looked at to see if we had a similar issue. Albert conjectures that the photoelectrons on the "off" G2 or G9 are showing up as current on the G8 DIB.
We discussed leaving the D3 and D8 front segment thresholds in place and accepting a ~90% livetime during the light leak issue. Next tohban should keep an eye on the light leak issue
Attenuator Thresholds
Attenuator State HEX value deadtime Min A0 0x00 0.00 % Max A0 0x10 6.27 % Min A1 0x07 2.74 % Max A1 0x15 8.24 % Min A2 0x05 1.96 % Max A2 0x29 16.08 % Min A3 0x09 3.53 % Max A3 0xFF 100.00 %
Spacecraft Management
Decimation | Normal/Vigorous |
HLAT Decimation | Rear decimation weight 6, no front decimation |
Night time data (fronts) | +/- 4 minutes |
Night time data (rears) | +/- 4 minutes |
Require extra passes? | No |
Requirement for moving pointer? | No |
Attenuator operation | Working as expected |
Detector problems? | See notes above. |